# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 3.2 Elliptic stable envelopes
[2] Let ${\boldsymbol{\lambda}}^{\prime}$ and ${\boldsymbol{\lambda}}^{\prime\prime}$ be $r^{\prime}$-tuple and $r^{\prime\prime}$-tuple partitions with $|{\boldsymbol{\lambda}}^{\prime}|=n^{\prime},|{\boldsymbol{\lambda}}^{\prime\prime}|=n^{\prime\prime}$, respectively. Consider the elliptic stable envelopes $\mathrm{Stab}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z})$ and $\mathrm{Stab}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})$ for $\mathrm{E}_{T}({\cal M}(n^{\prime},r^{\prime}))$ and $\mathrm{E}_{T}({\cal M}(n^{\prime\prime},r^{\prime\prime}))$ with the equivariant parameters ${\bf u}^{\prime}=(u_{1}^{\prime},\cdots,u_{r^{\prime}}^{\prime})$, ${\bf u}^{\prime\prime}=(u^{\prime\prime}_{1},\cdots,u^{\prime\prime}_{r^{\prime\prime}})$, respectively. Let $\mathfrak{C}^{\prime}$ and $\mathfrak{C}^{\prime\prime}$ be chambers $|u^{\prime}_{1}|\ll\cdots\ll|u^{\prime}_{r^{\prime}}|$ and $|u^{\prime\prime}_{1}|\ll\cdots\ll|u^{\prime\prime}_{r^{\prime\prime}}|$, respectively with a common stability condition $t_{1}/t_{2}>0$. Then the following shuffle product gives the elliptic stable envelope $\mathrm{Stab}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z})$ for $\mathrm{E}_{T}({\cal M}(n^{\prime}+n^{\prime\prime},r^{\prime}+r^{\prime\prime}))$ with the chamber $\mathfrak{C}$ given by $|u^{\prime}_{1}|\ll\cdots\ll|u^{\prime}_{r^{\prime}}|\ll|u^{\prime\prime}_{1}|\ll\cdots\ll|u^{\prime\prime}_{r^{\prime\prime}}|$ with the stability condition $t_{1}/t_{2}>0$.
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z}){}$
@@
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\prod_{i=1}^{r^{\prime}}\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}{\theta(\hbar u^{\prime}_{i}/x_{b}^{\prime\prime})}\prod_{j=1}^{r^{\prime\prime}}\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}\theta(x^{\prime}_{a}/u^{\prime\prime}_{j})\right.{}$
@@
×$\left.\hskip 113.81102pt\qquad\qquad\times\begin{matrix}&\cr&\cr\end{matrix}{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}\hbar^{-r^{\prime\prime}}){\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})\right).$ (3.6)
@@
Here $x_{a}^{\prime}\ (a\in{\boldsymbol{\lambda}}^{\prime})$ and $x_{b}^{\prime\prime}\ (b\in{\boldsymbol{\lambda}}^{\prime\prime})$ are the Chern roots for the tautological bundle on ${\cal M}(n^{\prime},r^{\prime})$ and ${\cal M}(n^{\prime\prime},r^{\prime\prime})$, respectively, and we set $\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}=\{x_{a}^{\prime}\}_{a\in{\boldsymbol{\lambda}}^{\prime}}\cup\{x_{b}^{\prime\prime}\}_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}$.
@@
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})=\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})(-)^{\varepsilon({\boldsymbol{\lambda}},r)}\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{\theta(t_{1}x_{a}/x_{b})}{\theta(x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\theta(x_{a}/u_{i}),$ (3.7)
@@
The difference between $\mathrm{Stab}$ and $\widehat{\mathrm{Stab}}$ is essentially the Thom class of $T^{1/2}=T^{1/2}{\cal M}(n,r)$ i.e.
@@
$\widehat{S}^{Ell}_{\boldsymbol{\lambda}}({\bf x},{\bf u})$ ::=:= $\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}\frac{\theta(x_{b}/x_{a})\theta(t_{2}x_{a}/x_{b})}{\theta(t_{1}x_{b}/x_{a})\theta(\hbar x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}>\rho_{r_{i}}}\left(-\frac{\theta(\hbar u_{i}/x_{a})}{\theta(x_{a}/u_{i})}\right){}$ (3.9)
×$\qquad\times\left.\prod_{a,b\in{\boldsymbol{\lambda}}\atop{c_{a}=c_{b}\atop h_{a}>h_{b}}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}\prod_{a,b\in{\boldsymbol{\lambda}}\atop{c_{a}+1=c_{b}\atop h_{a}<h_{b}}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}.\right.$
@@
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z}){}$
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta(x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(t_{1}x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\prod_{i=1}^{r^{\prime}}\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}\left(-\frac{\theta(\hbar u^{\prime}_{i}/x_{b}^{\prime\prime})}{\theta(x^{\prime\prime}_{b}/u^{\prime}_{i})}\right)\right.{}$
×$\hskip 85.35826pt\qquad\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}\hbar^{-r^{\prime\prime}})\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})\Biggr{)}.$ (3.10)
@@
In the following sections, we also use the elliptic stable envelopes for $\mathrm{E}_{T}({\cal M}(n,r))$ with the opposite polarization $T^{1/2}_{opp}=\hbar(T^{1/2})^{\vee}$, the elliptic nome $p^{*}$ and the Kähler parameter $\mathfrak{z}^{*-1}$:
${\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1}):={\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigl{|}_{T^{1/2}\mapsto T^{1/2}_{opp},\ p\mapsto p^{*},\ \mathfrak{z}\mapsto\mathfrak{z}^{*-1}}$ (3.11)
@@
${\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})=\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})(-)^{\varepsilon^{*}({\boldsymbol{\lambda}},r)}\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{\theta^{*}(t_{2}x_{a}/x_{b})}{\theta^{*}(\hbar x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\theta^{*}(\hbar u_{i}/x_{a}).{}$
@@
Note that for $T^{1/2}_{opp}=\hbar(T^{1/2}{\cal M}(n,r))^{\vee}$, we have
@@
$\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z}^{*-1}){}$
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta^{*}(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta^{*}(\hbar x^{\prime\prime}_{b}/x^{\prime}_{a})}{\theta^{*}(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta^{*}(t_{2}x^{\prime\prime}_{b}/x^{\prime}_{a})}\prod_{i=1}^{r^{\prime\prime}}\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}\left(-\frac{\theta^{*}(x^{\prime}_{a}/u^{\prime\prime}_{i})}{\theta^{*}(\hbar u^{\prime\prime}_{i}/x_{a}^{\prime})}\right)\right.{}$
×$\hskip 85.35826pt\qquad\qquad\times\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C}^{\prime\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z}^{*-1})\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C}^{\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}^{*-1}\hbar^{-r^{\prime\prime}})\Biggr{)}.$ (3.13)
Proof. The statement follows from (3.12) and the fact that ${\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ satisfies the same shuffle product formula as (3.6) except for replacing all $\theta$ with $\theta^{*}$. ∎
4 Vertex Operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
In this section, we introduce new vertex operators of type I and the type II dual by combining representations of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ in Sec.2.3 and the elliptic stable envelopes for $\mathrm{E}_{T}({\cal M}(n,r))$. We show that they yield correct shuffle product formula of the elliptic stable envelopes (Sec.4.3), the $\mathrm{K}$-theoretic vertex functions (Sec.5), exchange relations (Sec.6) and $L$-operator (Sec.7).
4.1 OPE of the elliptic currents
Let us consider the level $(1,N)$ representation given in Sec.2.3.1, on which $p^{*}=p\hbar^{-1}$. For the elliptic currents $x^{+}(u),x^{-}(v)$, one gets the following operator product expansion (OPE).

### v2

@@ 3.2. Elliptic stable envelopes
[2] Let $\boldsymbol{\lambda}$$^{\prime}$ and $\boldsymbol{\lambda}$$^{\prime\prime}$ be $r^{\prime}$-tuple and $r^{\prime\prime}$-tuple partitions with $|{\boldsymbol{\lambda}}^{\prime}|=n^{\prime},|{\boldsymbol{\lambda}}^{\prime\prime}|=n^{\prime\prime}$, respectively. Consider the elliptic stable envelopes $\mathrm{Stab}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z})$ and $\mathrm{Stab}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})$ for $\mathrm{E}_{T}({\mathcal{M}}(n^{\prime},r^{\prime}))$ and $\mathrm{E}_{T}({\mathcal{M}}(n^{\prime\prime},r^{\prime\prime}))$ with the equivariant parameters ${\bf u}^{\prime}=(u_{1}^{\prime},\cdots,u_{r^{\prime}}^{\prime})$, ${\bf u}^{\prime\prime}=(u^{\prime\prime}_{1},\cdots,u^{\prime\prime}_{r^{\prime\prime}})$, respectively. Let $\mathfrak{C}^{\prime}$ and $\mathfrak{C}^{\prime\prime}$ be chambers $|u^{\prime}_{1}|>\cdots>|u^{\prime}_{r^{\prime}}|$ and $|u^{\prime\prime}_{1}|>\cdots>|u^{\prime\prime}_{r^{\prime\prime}}|$, respectively with a common stability condition $t_{1}/t_{2}>0$. Then the following shuffle product gives the elliptic stable envelope $\mathrm{Stab}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z})$ for $\mathrm{E}_{T}({\mathcal{M}}(n^{\prime}+n^{\prime\prime},r^{\prime}+r^{\prime\prime}))$ with the chamber $\mathfrak{C}$ given by $|u^{\prime}_{1}|>\cdots\>|u^{\prime}_{r^{\prime}}|>|u^{\prime\prime}_{1}|>\cdots>|u^{\prime\prime}_{r^{\prime\prime}}|$ with the stability condition $t_{1}/t_{2}>0$.
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z})$ (3.5)
@@
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\right$.{}
@@
$\hskip 85.35826pt\times\begin{matrix}&\cr&\cr\end{matrix}\prod_{i=1}^{r^{\prime}}\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}{\theta(\hbar u^{\prime}_{i}/x_{b}^{\prime\prime})}\prod_{j=1}^{r^{\prime\prime}}\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}\theta(x^{\prime}_{a}/u^{\prime\prime}_{j}){}$
$\left.\hskip 71.13188pt\times\begin{matrix}&\cr&\cr&\cr\end{matrix}{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}\hbar^{-r^{\prime\prime}}){\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})\right).{}$
@@
Here $x_{a}^{\prime}\ (a\in{\boldsymbol{\lambda}}^{\prime})$ and $x_{b}^{\prime\prime}\ (b\in{\boldsymbol{\lambda}}^{\prime\prime})$ are the Chern roots for the tautological bundle on ${\mathcal{M}}(n^{\prime},r^{\prime})$ and ${\mathcal{M}}(n^{\prime\prime},r^{\prime\prime})$, respectively, and we set $\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}=\{x_{a}^{\prime}\}_{a\in{\boldsymbol{\lambda}}^{\prime}}\cup\{x_{b}^{\prime\prime}\}_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}$.
@@
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})$ (3.6)
$=\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})(-)^{\varepsilon({\boldsymbol{\lambda}},r)}\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{\theta(t_{1}x_{a}/x_{b})}{\theta(x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\theta(x_{a}/u_{i}),{}$
@@
The difference between $\mathrm{Stab}$ and $\widehat{\mathrm{Stab}}$ is essentially the Thom class of $T^{1/2}=T^{1/2}{\mathcal{M}}(n,r)$ i.e.
@@
$\widehat{S}^{Ell}_{\boldsymbol{\lambda}}({\bf x},{\bf u})$ (3.8)
$:=\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}\frac{\theta(x_{b}/x_{a})\theta(t_{2}x_{a}/x_{b})}{\theta(t_{1}x_{b}/x_{a})\theta(\hbar x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}>\rho_{r_{i}}}\left(-\frac{\theta(\hbar u_{i}/x_{a})}{\theta(x_{a}/u_{i})}\right){}$
$\qquad\qquad\times\left.\prod_{a,b\in{\boldsymbol{\lambda}}\atop{c_{a}=c_{b}\atop h_{a}>h_{b}}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}\prod_{a,b\in{\boldsymbol{\lambda}}\atop{c_{a}+1=c_{b}\atop h_{a}<h_{b}}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}.\right$.{}
@@
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z})$ (3.9)
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta(x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(t_{1}x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\right$.{}
$\hskip 71.13188pt\qquad\times\prod_{i=1}^{r^{\prime}}\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}\left(-\frac{\theta(\hbar u^{\prime}_{i}/x_{b}^{\prime\prime})}{\theta(x^{\prime\prime}_{b}/u^{\prime}_{i})}\right){}$
$\hskip 71.13188pt\quad\times\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}\hbar^{-r^{\prime\prime}})\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z})\Biggr{)}.{}$
@@
In the following sections, we also use the elliptic stable envelopes for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$ with the opposite polarization $T^{1/2}_{opp}=\hbar(T^{1/2})^{\vee}$, the elliptic nome $p^{*}$ and the Kähler parameter $\mathfrak{z}^{*-1}$:
${\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1}):={\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigl{|}_{p\mapsto p^{*},\ \mathfrak{z}\mapsto\mathfrak{z}^{*-1}}$ (3.10)
@@
$\hskip 8.53581pt{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ (3.11)
$\hskip-28.45274pt=\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})(-)^{\varepsilon^{*}({\boldsymbol{\lambda}},r)}\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{\theta^{*}(t_{2}x_{a}/x_{b})}{\theta^{*}(\hbar x_{a}/x_{b})}\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\theta^{*}(\hbar u_{i}/x_{a}).{}$
@@
Note that for $T^{1/2}_{opp}=\hbar(T^{1/2}{\mathcal{M}}(n,r))^{\vee}$, we have
@@
$\hskip 5.69054pt\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}(({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime});\mathfrak{z}^{*-1})$ (3.12)
$=\mathrm{Sym}_{\{x_{a}\}_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}}\left(\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}\frac{\theta^{*}(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta^{*}(\hbar x^{\prime\prime}_{b}/x^{\prime}_{a})}{\theta^{*}(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta^{*}(t_{2}x^{\prime\prime}_{b}/x^{\prime}_{a})}\right$.{}
$\hskip 56.9055pt\qquad\qquad\times\prod_{i=1}^{r^{\prime\prime}}\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}\left(-\frac{\theta^{*}(x^{\prime}_{a}/u^{\prime\prime}_{i})}{\theta^{*}(\hbar u^{\prime\prime}_{i}/x_{a}^{\prime})}\right){}$
$\hskip 2.84544pt\qquad\qquad\times\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C}^{\prime\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime\prime};\mathfrak{z}^{*-1})\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C}^{\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}^{*-1}\hbar^{-r^{\prime\prime}})\Biggr{)}.{}$
Proof. The statement follows from (3.11) and the fact that
${\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ satisfies the same shuffle product formula as (3.5) except for replacing all $\theta$ with $\theta^{*}$. ∎
4. Vertex Operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
In this section, we introduce new vertex operators, the type I and the type II dual, by combining representations of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ in Sec.2.3, 2.4 and the elliptic stable envelopes for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$. We show that they yield correct shuffle product formula of the elliptic stable envelopes (Sec.4.3), the $\mathrm{K}$-theoretic vertex functions (Sec.5), exchange relations (Sec.6) and $L$-operator (Sec.7).
4.1. OPEs of the elliptic currents
Let us consider the level $(1,N)$ representation given in Sec.2.3, on which $p^{*}=p\hbar^{-1}$. For the elliptic currents $x^{+}(u),x^{-}(v)$, one gets the following operator product expansion (OPE).